OpenStudy (anonymous):

Part 1 [3 points]: Find the solution to the equation 64^(3 + x) = 4^(2x). Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation.

OpenStudy (anonymous):

Find a common base between 4 and 64. Since 4^3 is 64, change the 64 to 4^3 so that it is now 4^3(4-x). Now that the bases are the same, just solve the equation 3(4-x)=2x. 4-x=6x 4=7x 4/7=x

OpenStudy (anonymous):

thanks :)